A mixture of H2(g), O2(g) and 2 mL of H2O (l) is present in a 0.500 L rigid container at 25 deg C. The number of moles of H2 and the number of moles of O2 are equal. The total pressure is 1146 mmHg and the vapor pressure of pure water is 24 mmHg. For all questions use the ideal gas constant (R) of 62.4 (L x mmHg)/(K x mol)

a. Determine the partial pressure of H2(g) and the partial pressure of O2(g) in the container.
b. If the reaction occurs at 25 deg C, how many grams of each gas, Hs(g) and O2(g) are present?
c. If the gas were moved to a balloon and the pressure remained constant, what volume would the gas occupy at 15 deg C?
d. If the liquid water is removed and the container is cooled, would the pressure in the container increase, decrease or remain the same. Explain
1
Expert's answer
2020-03-18T12:57:14-0400

(a)Total pressure is "1146mmHg"

Vapour Pressure of pure water is "24 mmHg"

Pressure by O2 and H2 is "1122 mmHg"

Mole fraction of H2 is "\frac{1}{2}" .

"P_{H_2}=\frac{1}{2}×1122=561mmHg."

"P_{O_2}=\frac{1}{2}×1122=561mmHg" .

(b)"PV=nRT"

"561\u00d70.5=n\u00d762.4\u00d7298"

"n=0.0151moles"

"M_{O_2}=0.0151\u00d732=0.483gm"

"M_{H_2}=0.0151\u00d72=0.0302gm"

(c)"PV=nRT"

"1122\u00d7V=0.0151\u00d72\u00d762.4\u00d7288"

V=0.484L

(d)"P\/T=constant"

"P=kT"

If temperature is decreased then pressure decreases.